# SPARKE: Scalable Prompt-Aware Diversity Guidance in Diffusion Models via RKE Score

### v1

@@ 4 Scalable Prompt-Aware Diversity Guidance in Diffusion Models
Utilizing the kernel-based entropy diversity scores and latent diffusion models (LDMs) described in Section 3.2, we introduce a scalable conditional entropy-based framework extending the recent Vendi score-based approach [10] in diversity-guided generative modeling to prompt-aware diversity enhancement that suits the conditional text-to-image models. Specifically, we first propose extending the order-1 Vendi score to the general matrix-based order-$\alpha$ Rényi entropy and subsequently to the conditional order-$\alpha$ Rényi entropy for achieving higher prompt-conditioned diversity. In addition, we demonstrate a computationally efficient special case in this family of diversity score functions by considering the order-2 entropy measures. This choice of entropy function leads to the RKE-guidance and Conditional-RKE guidance approaches for enhancing overall diversity and prompt-aware diversity in sample generation via diffusion models.
@@ 4.1 Diversity guidance via general order-$\alpha$ Renyi Entropy and RKE scores
The existing diversity score-based guidance approach, including the Vendi guidance in [10], optimizes the diversity score of the embedded version of the generated outputs, e.g. the CLIP embedding [52] of image data generated by the prompt-guided diffusion models. However, such a guidance process by considering an embedding on top of the diffusion model is computationally expensive and may not lead to semantically diverse samples. As discussed in [13, 42], computing the order-1 Vendi score and its gradients requires at least $\Omega(n^{2.367})$ computations for $n$ samples, and in practice involves $O(n^{3})$ computations for the eigen-decompositions of kernel matrices. This computational complexity limits practical window sizes of the existing Vendi guidance approach to a few hundred samples, i.e. the guidance function cannot be computed by a standard GPU processor for a sample size $n>500$.
To reduce computational complexity, we extend the order-1 Vendi score guidance to the general order-$\alpha$ Rènyi kernel entropy, and adopt the specific case $\alpha=2$, referred to as the RKE score. This formulation significantly lowers computational cost and is defined as:
$\mathcal{L}_{\text{RKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})=\frac{1}{\|\widetilde{K}_{Z}\|_{F}^{2}}=\frac{\mathrm{Tr}(K_{Z})^{2}}{\sum_{i=1}^{n}\sum_{j=1}^{n}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}},\quad\text{where}\,\,\widetilde{K}_{Z}=\frac{K_{Z}}{\mathrm{Tr}(K_{Z})},$ (6)
with kernel matrix $(K_{Z})_{ij}=k(\bm{z}^{(i)},\bm{z}^{(j)})$ computed directly on latent representations and $n$ denotes the total number of previously generated samples. $\mathcal{L}_{\text{RKE}}$ reduces the computational complexity from $O(n^{3})$ in order-1 Vendi score to $O(n^{2})$, enabling significantly larger batch sizes during the guided sampling process and leading to greater computational efficiency.
Input: Latents $\{\bm{z}^{(1)},\ldots,\bm{z}^{(n-1)}\}$, prompt features $\{\bm{y}^{(1)},\ldots,\bm{y}^{(n-1)}\}$, kernel functions $k_{Z}$, $k_{Y}$, denoising model $\epsilon_{\theta}$, diffusion reverse sampler, guidance scale $w$, diversity guidance scale $\eta$, decoder $\mathcal{D}$
Output: Diverse generated samples $\{\tilde{\bm{x}}^{(n)}\}$ with prompt features $\bm{y}^{(n)}$
1 for $t=T$ to $1$ do
2       Compute CFG guided noise: $\hat{\bm{\epsilon}}_{\theta}^{(n)}=(1+w)\cdot\bm{\epsilon}_{\theta}(\bm{z}_{t}^{(n)},t,\bm{y}^{(n)})-w\cdot\bm{\epsilon}_{\theta}(\bm{z}_{t}^{(n)},t)$ ;
3       Perform one-step denoising: $\bm{z}_{t-1}^{(n)}\leftarrow\text{Sampler}(\bm{z}_{t}^{(n)},\hat{\bm{\epsilon}}_{\theta}^{(n)})$ ;
4       Compute $K_{Z}$, $K_{Y}$, and $\mathcal{L}_{\text{Cond-IRKE}}=\|\widetilde{K}_{Z}\odot\widetilde{K}_{Y}\|_{F}^{2}$ ;
5       Compute the gradients $g^{(n)}=\nabla_{\bm{z}^{(n-1)}}\mathcal{L}_{\text{Cond-IRKE}}$ ;
6       Update latents: $\bm{z}_{t-1}^{(n)}\leftarrow\bm{z}_{t-1}^{(n)}-\eta\cdot g^{(n)}$ ;
7
Decode final samples: $\tilde{\bm{x}}^{(n)}=\mathcal{D}(\bm{z}_{0}^{(n)})$
Algorithm 1 Scalable Prompt-Aware Latent IRKE Diversity Guidance (SPARKE)
@@ 4.2 Diversity-Guided Sampling in Latent Diffusion Models
In contrast to previous work using order-1 Vendi score as the diversity guidance in ambient image space [10], we integrate efficient IRKE score diversity guidance directly into latent-space sampling. Standard LDMs employ classifier-free guidance (CFG) [46] for conditional sampling, iteratively denoising latents $\bm{z}_{t-1}$ from the noisy latents $\bm{z}_{t}$ at time step $t$ via reverse samplers:
$\bm{z}_{t-1}\leftarrow\text{Sampler}(\bm{z}_{t},\hat{\bm{\epsilon}}_{\theta}(\bm{z}_{t},t,\bm{y})),$ (7)
@@
Let $Z=\{\bm{z}^{(1)},\bm{z}^{(2)},\ldots,\bm{z}^{(n)}\}$ denote a set of $n$ generated data. Let kernel function $k:\mathcal{Z}\times\mathcal{Z}\rightarrow\mathbb{R}$ be symmetric and normalized, i.e. $k(\bm{z},\bm{z})=1$ for every $\bm{z}\in\mathcal{Z}$. Then, we observe that the function $\mathcal{L}_{\text{RKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})$ in Eq. equation 6 changes montonically with the Inverse-RKE function $\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})=1/\mathcal{L}_{\text{RKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})$, whose gradient with respect to $\bm{z}^{(n)}$ is:
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{IRKE}}\propto\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})\nabla_{\bm{z}^{(n)}}k(\bm{z}^{(i)},\bm{z}^{(n)}).$ (8)
The above Proposition 1 implies that an optimization objective maximizing $\mathcal{L}_{\text{RKE}}$ to promote diversity can be equivalently pursued by minimizing $\mathcal{L}_{\text{IRKE}}$, reducing the computational complexity from $O(n^{2})$ to $O(n)$. Leveraging this computational efficacy, our approach directly updates the gradient of $\mathcal{L}_{\text{IRKE}}$ rather than computing the gradient of $\mathcal{L}_{\text{RKE}}$ to the latent $\bm{z}^{(n)}$, allowing for an efficient, explicit diversity-promoting update. Specifically, for Latent Diffusion Models (LDMs), we introduce the following diversity-guided sampling via Eq. equation 8 in each time step $t$:
$\bm{z}_{t}\leftarrow\bm{z}_{t}-\eta\cdot\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{IRKE}}\,\,\quad(\textbf{IRKE Diversity Guidance})$ (9)
@@ 4.3 Conditional Prompt-Aware Diversity Guidance
To explicitly capture prompt-conditioned diversity and focusing on computational efficiency of order-2, we introduce the Conditional RKE loss:
$\mathcal{L}_{\text{Cond-RKE}}(\bm{z}^{(1)},\ldots,\bm{z}^{(n)};\bm{y}^{(1)},\ldots,\bm{y}^{(n)}):=\frac{\bigl{\|}\widetilde{K}_{Y}\bigr{\|}^{2}_{F}}{\bigl{\|}\frac{{K}_{Y}\odot{K}_{Z}}{\mathrm{Tr}({K}_{Y}\odot{K}_{Z})}\bigr{\|}^{2}_{F}},\quad\text{where}\,\,\widetilde{K}_{Y}=\frac{K_{Y}}{\mathrm{Tr}(K_{Y})},$ (10)
where $(K_{Y})_{ij}=k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})$ denotes the similarity among condition (e.g., prompts) embeddings. Similarly, we reduce the computational complexity from $O(n^{2})$ to $O(n)$, by computing the gradient of Conditional Inverse-RKE (Cond-IRKE) defined in Proposition 2:
@@
Let $Z$ denote a set of $n$ generated data, and $Y=\{\bm{y}^{(1)},\bm{y}^{(2)},\ldots,\bm{y}^{(n)}\}$ is the set of corresponding conditions. Let the kernel function $k$ be symmetric and normalized. Then, we observe that the function $\mathcal{L}_{\text{Cond-RKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)};\bm{y}^{(1)},\ldots,\bm{y}^{(n)})$ in Eq. equation 10 changes montonically with the Conditional Inverse-RKE function $\mathcal{L}_{\text{Cond-IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)};\bm{y}^{(1)},\ldots,\bm{y}^{(n)})=\|\widetilde{K}_{Z}\odot\widetilde{K}_{Y}\|_{F}^{2}$, whose gradient with respect to $\bm{z}^{(n)}$ is:
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}\propto\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\nabla_{\bm{z}^{(n)}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)}).$ (11)
Then we propose the following prompt-aware diversity-guided sampling via Eq. equation 11 in each time step $t$:
[…]
Algorithm 1 provides an explicit outline of this procedure. This Cond-IRKE guidance explicitly aligns the diversity of generated samples with respective prompts, while significantly maintaining the computational efficiency in gradient calculation.
@@ 5 Numerical Results
Figure 3: Comparison of SPARKE (Conditional RKE) Guidance with baselines on 2D Gaussian distributions
@@
Baselines. We compare our method with CADS [14], Particle Guidance [49], and Contextualized Vendi Score Guidance (c-VSG) [10]. In experiments where the reference is not available we use the VSG version of [10].
Models. We used Stable Diffusion [53] and the larger-scale Stable Diffusion XL (SDXL) [5] in our experiments on text-to-image generation, and also provide additional experimental results on different SOTAs in the Appendix. We also used the ControlNet model [27] for pose-to-image generation tasks.
Evaluation. We compared our method with baselines in terms of output diversity and fidelity. We measured fidelity via IS [54], CLIPScore [55] and KID [56] for evaluating the quality and consistency of prompts. We also measured Density [37] in cases where a proper reference dataset was available. For evaluating diversity, we measured Coverage [37] given a reference dataset, and used the Vendi score [13] and Conditional-Vendi score [15] in cases with no proper reference dataset. To measure the diversity within each cluster of prompts, we used the in-batch similarity score introduced in [49] as a golden metric, which represents the average pairwise cosine similarity of features within an image batch. We note that this score can be evaluated if we know the ground-truth cluster of prompts. In addition to that, we report Authenticity (AuthPct) [57] to measure the generalization. As recommended in [58], we used the DINOv2 [59] backbone embedding for computing the evaluation scores.
Synthetic Datasets. To compare our method with previous works, we trained a diffusion model using the DDIM sampler [60] on 2D Gaussian mixture model benchmarks similar to [61]. We fine-tuned the hyperparameters as explained in detail in the Appendix for the tested methods. As shown in Figure 3, CADS was able to cover the modes’ support set, but exhibited a slight distribution shift, while c-VSG adhered better to the modes yet led to visually lower diversity. Particle Guidance (PG) and SPARKE covered the modes satisfactorily, with SPARKE generating fewer noisy samples due to its clustering-based structure compared to the sample-based comparison in PG.
Figure 4: Comparison of latent entropy guidance vs. CLIP-(ambient) entropy guidance on SD-XL.
[…]
(CLIP-embedded) Ambient VSG [10] 5.45 5.34 13.33 0.28
Latent VSG 7.24 4.42 25.22 0.15
Comparison of Entropy Guidance Effects in Latent and Ambient Spaces. We compared latent-space guidance (as in SPARKE) with the CLIP-embedded ambient-space guidance in c-VSG [10] using three categories from the GeoDE dataset [62] and modified it with GPT-4o [63] as explained in the Appendix. As qualitatively shown in Figure 4, CLIP-ambient led to visual artifacts and offered less improvement in output visual diversity, while latent guidance produced more semantically diverse outputs. Table 1 shows that although CLIP-based Vendi scores were slightly higher for CLIP-embedded ambient guidance, which could be due to optimizing Vendi with the CLIP features, the DINOv2-based Vendi scores and in-batch similarity scores indicate higher diversity for the latent Vendi entropy guidance that we proposed for LDMs. Furthermore, latent guidance can be performed with a considerably lower GPU memory (in our implementation $\approx$20GB for the latent case vs. $\approx$ 35GB for the ambient case), resulting in higher computational efficiency.
Effect of Classifier-Free Guidance scale on SPARKE guidance. We study the Classifier-free guidance’s impact on SPARKE for the Stable Diffusion 1.5 model, demonstrating how our method alleviates the known quality-diversity trade-off [46]. We changed the CFG scale from 2 to 8 while using the same setting of SD 1.5 as mentioned in Section C.2. We evaluated SPARKE’s effect on the quality-diversity trade-off using Density/Coverage metrics in Figure 6. Our results show that SPARKE guidance maintains diversity in a high CFG scale while preserving acceptable generation quality. Additionally, the Vendi score and Kernel Distance (KD) evaluations show the same trend. Qualitative comparisons of generated samples at CFG scales $w=4,6,8$ are presented in Figure 7, demonstrating that SPARKE produces diverse outputs without compromising quality.
[…]
Comparison of prompt-aware diversity guidance vs. existing unconditional guidance. We compared the prompt-aware diversity guidance with the existing prompt-unconditional diversity guidance methods. We considered the first 10,000 prompts from the MS-COCO 2014 validation set [64]. Applying Stable Diffusion 2.1, we generated samples using both latent RKE conditional guidance for prompt-aware diversity and an unconditional RKE guidance baseline targeting prompt-unaware diversity over all the generated images. The results in Figure 5 suggest that the diversity improvement in the unconditional guidance becomes less effective after more iterations pass. On the other hand, the prompt-aware diversity guidance in SPARKE remained effective over all 40k iterations of the experiment. Table 2 contains the quantitative evaluation results using SD-2.1, where prompt-aware methods achieved higher diversity while preserving image quality and prompt alignment. Additional results are provided in the Appendix D.
@@
SD (No Guidance) [53] 31.20 52.37 86.94 26.54 369.37 80.36
c-VSG [10] 28.75 56.25 87.91 27.91 376.48 79.82
CADS [14] 29.89 55.08 88.59 28.73 380.08 79.44
latent RKE Guidance (Ours) 30.18 55.37 88.26 29.88 387.59 79.01
SPARKE: latent Conditional-RKE (Ours) 30.96 53.15 89.10 32.57 405.51 75.68
6 Conclusion

### v2

@@ 4 Scalable Prompt-Aware Diversity Guidance in Diffusion Models
Utilizing the kernel-based entropy diversity scores and latent diffusion models (LDMs) described in Section 3.2, we introduce a scalable conditional entropy-based framework extending the recent Vendi score-based approach askari2024improving in diversity-guided generative modeling to prompt-aware diversity enhancement that suits the conditional text-to-image models. Specifically, we first propose extending the order-1 Vendi score to the general matrix-based order-$\alpha$ Rényi entropy and subsequently to the conditional order-$\alpha$ Rényi entropy for achieving higher prompt-conditioned diversity. In addition, we demonstrate a computationally efficient special case in this family of diversity score functions by considering the order-2 entropy measures. This choice of entropy function leads to the RKE-guidance and Conditional-RKE guidance approaches for enhancing overall diversity and prompt-aware diversity in sample generation via diffusion models.
@@ 4.1 Diversity guidance via general order-$\alpha$ Renyi Entropy and RKE scores
The existing diversity score-based guidance approach, including the Vendi guidance in askari2024improving , optimizes the diversity score of the embedded version of the generated outputs, e.g. the CLIP embedding radford2021learning of image data generated by the prompt-guided diffusion models. However, such a guidance process by considering an embedding on top of the diffusion model is computationally expensive and may not lead to semantically diverse samples. As discussed in friedman2023vendi ; ospanov2024towards , computing the order-1 Vendi score and its gradients requires at least $\Omega(n^{2.367})$ computations for $n$ samples, and in practice involves $O(n^{3})$ computations for the eigen-decompositions of kernel matrices. This computational complexity limits practical window sizes of the existing Vendi guidance approach to a few hundred samples, i.e. the guidance function cannot be computed by a standard GPU processor for a sample size $n>500$.
To reduce computational complexity, we extend the order-1 Vendi score guidance to the general order-$\alpha$ Rènyi kernel entropy, and adopt the specific case $\alpha=2$, referred to as the RKE score jalali2023information . This formulation significantly lowers computational cost and is defined as:
$\mathcal{L}_{\text{RKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})=\Bigl\|\tfrac{1}{n}K_{Z}\Bigr\|_{F}^{-2}=\Bigl(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k^{2}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})\Bigr)^{-1},$ (5)
with kernel matrix $(K_{Z})_{ij}=k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})$ computed directly for the latent representations and $n$ denotes the total number of generated samples. Therefore, applying $\mathcal{L}_{\text{RKE}}$ as the guidance potential function reduces the computational complexity from $O(n^{3})$ in order-1 Vendi score to $O(n^{2})$, enabling significantly larger batch sizes during the guided sampling process and resulting in higher efficiency.
@@ 4.2 Diversity-Guided Sampling in Latent Diffusion Models
In contrast to previous work using order-1 Vendi score as the diversity guidance in ambient image space askari2024improving , we integrate efficient IRKE score diversity guidance directly into latent-space sampling. Standard LDMs employ classifier-free guidance (CFG) ho2022classifier for conditional sampling, iteratively denoising latents $\boldsymbol{z}_{t-1}$ from the noisy latents $\boldsymbol{z}_{t}$ at time step $t$ via reverse samplers:
$\boldsymbol{z}_{t-1}\leftarrow\text{Sampler}(\boldsymbol{z}_{t},\hat{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{z}_{t},t,\boldsymbol{y})),$ (6)
@@
Let $Z=\{\boldsymbol{z}^{(1)},\boldsymbol{z}^{(2)},\ldots,\boldsymbol{z}^{(n)}\}$ denote a set of $n$ generated data. Let kernel function $k:\mathcal{Z}\times\mathcal{Z}\rightarrow\mathbb{R}$ be symmetric and normalized, i.e. $k(\boldsymbol{z},\boldsymbol{z})=1$ for every $\boldsymbol{z}\in\mathcal{Z}$. Then, we observe that the function $\mathcal{L}_{\text{RKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})$ in Eq. (5) changes montonically with the Inverse-RKE function $\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})=1/\mathcal{L}_{\text{RKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})$, whose gradient with respect to $\boldsymbol{z}^{(n)}$ is:
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{IRKE}}\propto\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\nabla_{\boldsymbol{z}^{(n)}}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}).$ (7)
The above Proposition 1 implies that an optimization objective maximizing $\mathcal{L}_{\text{RKE}}$ to promote diversity can be equivalently pursued by minimizing $\mathcal{L}_{\text{IRKE}}$, reducing the computational complexity from $O(n^{2})$ to $O(n)$. Leveraging this computational efficacy, our approach directly updates the gradient of $\mathcal{L}_{\text{IRKE}}$ rather than computing the gradient of $\mathcal{L}_{\text{RKE}}$ to the latent $\boldsymbol{z}^{(n)}$, allowing for an efficient, explicit diversity-promoting update. Specifically, for Latent Diffusion Models (LDMs), we introduce the following diversity-guided sampling via Eq. (7) in each time step $t$:
$\boldsymbol{z}_{t}\leftarrow\boldsymbol{z}_{t}-\eta\cdot\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{IRKE}}\,\,\quad(\textbf{IRKE Diversity Guidance})$ (8)
@@
Input: Latents $\{\boldsymbol{z}^{(1)},\ldots,\boldsymbol{z}^{(n-1)}\}$, prompt features $\{\boldsymbol{y}^{(1)},\ldots,\boldsymbol{y}^{(n-1)}\}$, kernel functions $k_{Z}$, $k_{Y}$, denoising model $\epsilon_{\theta}$, diffusion reverse sampler, guidance scale $w$, diversity guidance scale $\eta$, decoder $\mathcal{D}$
Output: Diverse generated samples $\{\tilde{\boldsymbol{x}}^{(n)}\}$ with prompt features $\boldsymbol{y}^{(n)}$
1 for $t=T$ to 11 do
2   Compute CFG guided noise: $\hat{\boldsymbol{\epsilon}}_{\theta}^{(n)}=(1+w)\cdot\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t}^{(n)},t,\boldsymbol{y}^{(n)})-w\cdot\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t}^{(n)},t)$ ;
3   Perform one-step denoising: $\boldsymbol{z}_{t-1}^{(n)}\leftarrow\text{Sampler}(\boldsymbol{z}_{t}^{(n)},\hat{\boldsymbol{\epsilon}}_{\theta}^{(n)})$ ;
4   Compute $K_{Z}$, $K_{Y}$, and $\mathcal{L}_{\text{Cond-IRKE}}=\frac{1}{n^{2}}\bigl\|{K}_{Z}\odot{K}_{Y}\bigr\|_{F}^{2}$ ;
5   Compute the gradients $g^{(n)}=\nabla_{\boldsymbol{z}^{(n-1)}}\mathcal{L}_{\text{Cond-IRKE}}$ ;
6   Update latents: $\boldsymbol{z}_{t-1}^{(n)}\leftarrow\boldsymbol{z}_{t-1}^{(n)}-\eta\cdot g^{(n)}$ ;
7
Decode final samples: $\tilde{\boldsymbol{x}}^{(n)}=\mathcal{D}(\boldsymbol{z}_{0}^{(n)})$
Algorithm 1 Scalable Prompt-Aware Latent IRKE Diversity Guidance
@@ 4.3 Conditional Prompt-Aware Diversity Guidance
To explicitly capture prompt-conditioned diversity and focus on the computational efficiency of order-2, we introduce the Conditional RKE loss:
$\mathcal{L}_{\text{Cond-RKE}}(\boldsymbol{z}^{(1)},\ldots,\boldsymbol{z}^{(n)};\boldsymbol{y}^{(1)},\ldots,\boldsymbol{y}^{(n)}):=\frac{\bigl\|K_{Y}\bigr\|^{2}_{F}}{\bigl\|K_{Y}\odot K_{Z}\bigr\|^{2}_{F}},$ (9)
where $(K_{Y})_{ij}=k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})$ denotes the similarity among condition (e.g., prompts) embeddings. Similarly, we reduce the computational complexity from $O(n^{2})$ to $O(n)$, by computing the gradient of Conditional Inverse-RKE (Cond-IRKE) defined in Proposition 2:
@@
Let $Z$ denote a set of $n$ generated data, and $Y=\{\boldsymbol{y}^{(1)},\boldsymbol{y}^{(2)},\ldots,\boldsymbol{y}^{(n)}\}$ is the set of corresponding conditions. Let the kernel function $k$ be symmetric and normalized. Then, we observe that the function $\mathcal{L}_{\text{Cond-RKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)};\boldsymbol{y}^{(1)},\ldots,\boldsymbol{y}^{(n)})$ in Eq. (9) changes montonically with the Conditional Inverse-RKE function $\mathcal{L}_{\text{Cond-IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)};\boldsymbol{y}^{(1)},\ldots,\boldsymbol{y}^{(n)})=\|\widetilde{K}_{Z}\odot\widetilde{K}_{Y}\|_{F}^{2}$, whose gradient with respect to $\boldsymbol{z}^{(n)}$ is:
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}\propto\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\nabla_{\boldsymbol{z}^{(n)}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}).$ (10)
Then we propose the following prompt-aware diversity-guided sampling via Eq. (10) in each time step $t$:
$\boldsymbol{z}_{t}\leftarrow\boldsymbol{z}_{t}-\eta\cdot\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}\,\,\quad(\textbf{Cond-IRKE Diversity Guidance})$ (11)
Algorithm 1 provides an explicit outline of this procedure. This Cond-IRKE guidance explicitly aligns the diversity of generated samples with respective prompts, while significantly maintaining the computational efficiency in gradient calculation. In Appendix B.3, we provide a theoretical interpretation based on a stochastic differential equation.
@@ 5 Numerical Results
Figure 4: Comparison of SPARKE (Conditional RKE) Guidance with baselines on 2D GMMs.
@@
Baselines. We compare our method with CADS sadat2024cads , Particle Guidance corso2024particleguidance , and Contextualized Vendi Score Guidance (c-VSG) askari2024improving . In experiments without a reference, we use VSG askari2024improving .
Models. We used Stable Diffusion Rombach_2022_CVPR and the larger-scale Stable Diffusion XL (SDXL) podell2024sdxl , and PixArt-$\Sigma$ pixart_sigma in our experiments on text-to-image generation. We also provide additional experimental results on different SOTAs in the Appendix.
@@
SD (No Guidance) Rombach_2022_CVPR 31.20 52.37 26.54 369.37 80.36
c-VSG askari2024improving 28.75 56.25 27.91 376.48 79.82
CADS sadat2024cads 29.89 55.08 28.73 380.08 79.44
latent RKE Guidance (Ours) 30.18 55.37 29.88 387.59 79.01
SPARKE: latent Cond-RKE (Ours) 30.96 53.15 32.57 405.51 75.68
Evaluation. We compared our method with baselines in terms of output diversity and fidelity. Fidelity was measured using CLIPScore clipscore , and KID bińkowski2021demystifyingmmdgans to evaluate prompt quality and consistency, and Density naeem2020reliable where a reference dataset was available. For diversity, we used Vendi friedman2023vendi and Conditional-Vendi jalali2024conditional scores, and we used Coverage naeem2020reliable when a reference dataset existed. We measured diversity within each prompt cluster using the in-batch similarity score corso2024particleguidance , calculated as the average pairwise cosine similarity of image features in a batch.
Synthetic Datasets. We compare our method on 2D Gaussian mixture benchmarks gulrajani2017improved , training a diffusion model with the DDIM sampler song2020denoising and fine-tuning hyperparameters (see Appendix C). Figure 4 shows that CADS covered the modes’ support but with a slight distribution shift, while c-VSG better adhered to the modes at the cost of lower diversity. PG and SPARKE both covered the modes well, with SPARKE producing fewer noisy samples due to its clustering-based structure.
Comparison of Entropy Guidance Effects in Latent and Ambient Spaces. We compared latent-space guidance (SPARKE) with CLIP-embedded ambient-space guidance from c-VSG askari2024improving using modified GeoDE categories ramaswamy2022geode . As qualitatively shown in Figure 2, CLIP-ambient led to visual artifacts and offered less improvement in output visual diversity, while latent guidance produced more semantically diverse outputs. We also provided quantitative results in Table 5, which show that latent guidance achieves higher diversity. Also, latent guidance requires significantly less GPU memory ($\approx$20GB vs. $\approx$35GB), offering substantially better computational efficiency.
Comparison of Prompt-Aware Diversity Guidance vs. Unconditional Guidance. We generated images for 10,000 MS-COCO prompts mscoco using Stable Diffusion 2.1, comparing latent RKE conditional guidance (prompt-aware) against unconditional RKE guidance (prompt-unaware). Figure 5 shows unconditional diversity guidance becomes less effective over time, while SPARKE’s prompt-aware approach remained effective throughout all 40k iterations. Table 1 confirms that prompt-aware methods achieve higher diversity while preserving image quality and prompt alignment. Also, Table 2 provides the quantitative results for SD-XL and PixArt-$\Sigma$.
Table 2: Comparison of diversity and fidelity metrics across different models and methods.
[…]
SDXL 31.17 66.37 27.54 309.54 81.67
SDXL + SPARKE 30.47 62.62 31.17 313.80 77.75
PixArt-$\Sigma$ 31.01 65.37 26.45 307.36 83.84
PixArt-$\Sigma$ + SPARKE 30.66 63.86 32.14 322.25 78.81
Novelty Guidance Using SPARKE. We evaluated novelty guidance by using the SPARKE method to guide an SD-XL model with a reference dataset of cat and dog images. As shown in Figure 6, this approach successfully reduced the generation of cats and dogs, yielding more novel samples with respect to the reference dataset and resulting in a more balanced distribution of other animals.
Figure 6: Novelty guidance with SPARKE. Using the prompt "A young lady walking with an animal in the park," samples are generated with respect to a reference set of cat and dog images, resulting in more novel samples compared to the base model.
Effect of Classifier-Free Guidance scale on SPARKE guidance. We study the Classifier-free guidance’s impact on SPARKE for the Stable Diffusion 1.5 model, demonstrating how our method alleviates the known quality-diversity trade-off ho2022classifier . We changed the CFG scale from 2 to 8 and we evaluated SPARKE’s effect on the quality-diversity trade-off using Density/Coverage metrics in Figure 7. Our results show that SPARKE guidance maintains diversity in a high CFG scale while preserving acceptable generation quality. Additionally, the Vendi score and Kernel Distance (KD) evaluations show the same trend. Qualitative comparisons of generated samples at CFG scales $w=4,6,8$ are presented in Figure 8, demonstrating that SPARKE produces diverse outputs without compromising quality.
[…]
Figure 8: Qualitative comparison of DiT-XL-2 with different classifier-free guidance scales: (Left side) standard DiT-XL-2 (Right side) DiT-XL-2 with SPARKE prompt-aware diversity guidance.
Computational Efficiency Comparison. We evaluated SPARKE’s computational efficiency by comparing its runtime and peak memory with the baselines over 1000 text-to-image generations using a 50-step diffusion process on an NVIDIA RTX 4090. The results, as shown in Table 3, indicate that SPARKE requires significantly lower runtime and less memory (as it functions in the latent space).
Table 3: Comparison of runtime and GPU memory usage for different guidance methods.
Model + Guidance Method Runtime per sample (s) GPU Memory Peak (GB)
Stable Diffusion v1.5 1.620 $\pm$ 0.265 3.178 $\pm$ 0.001
SD-1.5 + SPARKE (Ours) 1.752 $\pm$ 0.287 3.230 $\pm$ 0.025
SD-1.5 + c-VSG 3.079 $\pm$ 0.296 8.665 $\pm$ 0.116
SD-1.5 + Particle Guide (Pixel) 4.912 $\pm$ 0.312 7.531 $\pm$ 0.032
SD-1.5 + Particle Guide (DINOv2) 9.102 $\pm$ 0.271 20.133 $\pm$ 0.038
6 Conclusion and Limitations